### v1

@@ II-D Automatic Labeling And Human Recheck
Through the application of hallucination ranking, we can identify continuations that are both articulately expressed and likely to contain hallucinations. To detect continuations with confirmed hallucinations, we propose an annotation scheme that utilizes keywords, which includes automatic labeling and subsequent human verification, as shown in Fig. 4.
Figure 4: The process of automatic labeling and human recheck.
@@
Automatic labeling. We utilize the keywords identified by GPT3.5-Turbo from the candidate continuations, similarly to the process used in the computation of kwPrec previously. These keywords act as the focal points for subsequent verification. Thereafter, we employ GPT4-0613 [20] to perform annotation on these keywords. GPT4-0613 evaluates the validity of the keywords in the continuations by conducting a cross-reference with the provided original news and provides explanations for any detected unreasonable keywords.
@@
Human recheck. We undertake a manual, one-to-one verification process by analyzing the annotated results and explanations provided by GPT4-0613 against the original news. This step is implemented to ensure the accuracy of the machine-generated annotations. In the end, instances verified as accurate by annotators comprise the final $\mathbb{UHG}$Eval dataset.
However, the keyword-based annotation scheme exhibits inherent limitations. Languages exhibit a dependency structure among words [21]. For instance, in the phrase “The rainbow is black,” the words “rainbow” and “black” exhibit interdependence. One could contend that “black” is incorrect, while another could maintain that “rainbow” is the erroneous term, given that “night” is typically described as black. To address the annotation challenges stemming from language dependency structures, we have adopted the Least Hallucination Principle. If a set of words can be selected, and their replacement with contextually appropriate words yields a semantically coherent sentence, then such a set of words is designated as a hallucinated word group. The words selected for annotation must meet the condition of comprising the minimal number of words in the group, as illustrated in Equation 1. In the equation, $\mathbf{W}$ is the set of keywords in a sentence, $\mathbf{w}$ is the hallucinated word group, $\text{correct}(\cdot)$ is the correction function that modifies hallucinated words to non-hallucinated words, and $\text{hallucinated}(\cdot)$ assesses whether a sentence composed of a set of keywords hallucinated.
In accordance with this principle, within the phrase “Journey to the West is an American novel and one of the Four Great Classics,” the word “American” would be marked for annotation, as altering this single keyword to “Chinese” dispels the hallucination throughout the sentence.
Additionally, we acknowledge that the task of hallucination annotation may become somewhat tedious. Consequently, annotators are integrated throughout the entire process, participating in discussions instead of solely evaluating the accuracy of machine annotations. This approach also yields benefits for our work. For example, an annotator with a journalism background offered valuable professional insights into pinpointing news-related hallucinations, emphasizing that fact increment is a critical aspect of news writing.
II-E Data Statics
Starting with 17,714 candidate hallucinated continuations, we curated a dataset of 5,141 hallucinated continuations, as detailed in the basic statistics in Table II. Additionally, we developed a conversion rate chart to depict the transition from candidate hallucinations to the final dataset, as depicted in Fig. 5. The conversion rate can be interpreted as the likelihood of hallucinations occurring across various categories. Our observations indicate a higher likelihood of hallucinations in number-intensive and general news, whereas this likelihood is reduced in knowledge-intensive and document-intensive news.
TABLE II: Dataset Basic Statics
DOC KNO NUM GEN
#news 1242 320 2431 1148
avg. #hallu. kw. 2.15 1.99 2.54 2.12
avg. #kw. 8.43 8.09 8.07 8.17
#hallu. kw. / #kw. 25.47% 24.61% 31.44% 26.00%
avg. len. contn. 46.77 48.36 44.47 45.97
avg. len. begin. 102.15 102.66 103.20 102.86
avg. len. refer. 634.17 618.90 624.47 632.47
@@ II-E Data Statics
  • Note: In the table, # denotes quantity, avg. denotes average, len. denotes length, contn. denotes hallucinated continuations, begin. denotes news beginnings, and refer. denotes reference information. The same as below.
Figure 5: Conversion rates from candidates to hallucinations.
By analyzing the hallucinated word cloud depicted in Fig. 6 for each news category, we can draw the following conclusions: Number-intensive news often includes numeric values that are challenging to remember, like 0.09% and 6:3, which pose difficulties for both LLMs and humans. General news encompasses a diverse vocabulary, featuring terms such as “social media” and “friendship,” which are often deemed less critical and thus challenging to incorporate into the training corpora of many LLMs. Knowledge-intensive news frequently features terms such as “according to incomplete statistics” and “key technology,” which are prevalent in technical literature. However, LLMs may not always use these terms appropriately. Document-intensive news often contains terms associated with official statements, such as “representation,” “president,” and “spokesperson.” This suggests that LLMs are susceptible to introducing unauthorized alterations to the content documents.
Figure 6: Word clouds of hallucinated keywords in different types of news
III Experiments
III-A Models
@@
Given that our dataset is tailored for the Chinese language generation domain, we selected eight widely-used Chinese LLMs and three foundational models from OpenAI, as detailed in Table III. These include eight base models: GPT Base, GLM Base, BLOOMZ Base, InternLM Base, Baichuan2 Base, Qwen Base, Aquila2 Base, and LLaMA2 Base.
TABLE III: Models Sorted by Release Date
Model Parm. Type Publisher Release
GPT3.5-Turbo [1] $*{}^{*}$ Chat OpenAI $*{}^{*}$
GPT4-0613 [20] NaN Chat OpenAI 2023.06
ChatGLM2 [12] 6B Chat Tsinghua 2023.06
Xinyu 7B Chat IAAR&Xinhua 2023.06
InternLM [15] 20B Chat ShLab 2023.07
Baichuan2 [13] 13B Chat Baichuan Inc. 2023.09
Baichuan2 [13] 53B Chat Baichuan Inc. 2023.09
Qwen [14] 14B Chat Alibaba 2023.09
Aquila2 [22] 34B Chat BAAI 2023.10
Xinyu2 70B Chat IAAR&Xinhua 2023.10
GPT4-1106222https://openai.com/blog/new-models-and-developer-products-announced-at-devday NaN Chat OpenAI 2023.11
  • Note: In the table, asterisk (*) denotes estimated value, NaN denotes no public data available, and 175B denotes 175 billion.
GPT represents a series of LLMs developed by OpenAI [20]. In this study, GPT3.5-Turbo, GPT4-0613, and GPT4-1106 are utilized. GLM constitutes a pre-training framework proposed by Tsinghua University [12], and the ChatGLM2-6B chat model is employed. BLOOMZ is a variant derived via multitask prompted fine-tuning (MTF) of the pre-trained BLOOM model [16], and following supplementary training, it is integrated into Xinyu-7B. InternLM serves as an open-source, lightweight training framework, with its development team releasing a spectrum of models utilizing this framework [15]; the InternLM-20B open-source chat model is utilized in the present work. Baichuan2 comprises a series of expansive, multilingual base language models [13], with both the open-source Baichuan2-7B chat model and the closed-source Baichuan2-53B model being employed in this investigation. Qwen encompasses a language model series characterized by distinct models with varying parameter counts [14], and the Qwen-14B open-source chat model is utilized in the current study. Aquila2 represents a language model series devised by BAAI, noted for surpassing comparable models in terms of performance [22], and the Aquila2-34B chat model is employed in this research. LLaMA2 constitutes a suite of pre-trained and fine-tuned LLMs, with scales ranging from 7 billion to 70 billion parameters [17]. Following additional training, LLaMA2-70B is incorporated into Xinyu2-70B.
III-B Evaluation Method
For the evaluation of hallucinations in LLMs, the task is decomposed into three principal dimensions: form, metric, and granularity. Form concerns the manner in which the model interacts with the evaluation dataset; metric refers to the precise computational approach utilized for performance assessment; and granularity signifies the depth of detail considered in the evaluation of hallucinations.
In terms of form, this encompasses human evaluation, discriminative evaluation, selective evaluation, and generative evaluation, among others. Human evaluation entails the direct application of human judgment to determine if the model’s output contains hallucinations, representing a critical evaluation form [23]. However, the drawbacks of this approach are evident: evaluating in excess of 5000 data points is tantamount to creating a new dataset, with the associated time and financial expenditures proving prohibitive.
Discriminative evaluation enables LLMs to respond with binary answers of “yes” or “no” [6, 24]. Specifically, this evaluation modality involves presenting the LLM under scrutiny with an initial text followed by a continuation that may or may not include hallucinations. The LLM is tasked with producing a verdict as to the presence of hallucinations. Owing to the efficacy of few-shot prompting, this evaluation paradigm is relatively uncomplicated for LLMs to administer, as it facilitates the elicitation of the requisite responses. However, this method depends solely on the LLM’s ability to draw upon the knowledge encoded within its parameters, necessitating the concurrent application of knowledge and reasoning, and thus requiring a robust foundational model capacity.
Similar to discriminative evaluation, selective evaluation allows LLMs to tackle multiple-choice questions by choosing between option A or B, as exemplified by PandaLM [25]. Specifically, in selective evaluation, the LLM under evaluation is presented with an initial text followed by two continuations: one that includes hallucinations and another that does not. The LLM’s objective is to identify which of the two is hallucinated. This assessment method offers the LLM more contextual information than discriminative evaluation, thereby alleviating the burden of fact-checking and lessening the dependence on retrieving facts from its parameters. Consequently, this reduces the level of difficulty for the LLM.
However, both discriminative and selective evaluations encounter a substantial challenge. They are predicated on the assumption that “LLMs’s capacity to produce reliable text is contingent upon their discernment between hallucinated and non-hallucinated content.” These methods do not simulate the evaluation of the model’s output for hallucinations. Consequently, generative evaluation is crucial as it directly evaluates the presence of hallucinations in the text generated by the LLM. Specifically, the LLM under evaluation is provided with an initial text and is then tasked with generating a continuation. Subsequently, various reference-based techniques are utilized to determine if the continuation includes hallucinations. However, the challenge arises from the fact that it is not feasible to automatically and accurately ascertain if newly generated text is hallucinated; if it were, annotated datasets would be redundant. In scenarios of unrestrained text generation, this issue becomes increasingly complex. This complexity stems from the fact that text generated without constraints may introduce a multitude of entities and facts absent in the reference material, complicating the verification of their accuracy. Despite these hurdles, generative evaluation continues to be a predominant strategy in Natural Language Generation (NLG) tasks [26].
In terms of metrics, these include classification metrics such as accuracy, precision, recall, and others, which are applicable to human evaluation, discriminative evaluation, and selective evaluation. Generative evaluation, on the other hand, encompasses both lexical and semantic metrics. Lexical metrics evaluate the extent of token overlap between the generated text and the reference information, including metrics such as BLEU [18], ROUGE [19], and the newly proposed kwPrec. Semantic metrics gauge the similarity in meaning between sentences, with examples including BERTScore [27], GPT-judge [28], and GPTScore [29], among others.
In terms of granularity, evaluations can be conducted at both the sentence and keyword levels. Owing to our annotation methodology, our dataset is marked at the keyword level to signify instances of hallucinations. This approach affords a broader spectrum of possibilities for configuring the evaluation task, enabling the evaluated model to address the presence of hallucinations at either the sentence level or keyword level.
III-C Evaluation Framework
@@
In order to accommodate different forms of evaluation methods, we have developed a of data-secure, easy-to-extend and easy-to-use evaluation framework, as illustrated in Fig. 7.
@@
The framework comprises four ascending layers: the dependency layer, the evaluator layer, the core layer, and the interface layer. The dependency layer delineates the requisite underlying modules for the evaluation framework, encompassing datasets, LLM hubs, and diverse metrics. Notably, all underlying modules are extensible; datasets may be supplanted with customized versions, LLMs sourced from APIs or platforms such as Hugging Face [^3] https://huggingface.co/models, and metrics tailored individually. The evaluator layer, constituting the second tier, centers on an abstract class, Evaluator, and its various implementations. Within this layer, three distinct types are implemented: GenerativeEvaluator, DiscriminativeEvaluator, and SelectiveEvaluator. Users may also engineer custom evaluators, contingent upon adherence to the interface specifications of the abstract class, necessitating merely three function overloads. The core layer, representing the third stratum, comprises two principal modules: experiment.py and analyst.py. The former module facilitates experiments involving multiple LLMs, evaluators, and processes, whereas the latter module is tasked with the statistical analysis of experimental outcomes. The interface layer, constituting the final tier, orchestrates the user’s interaction with $\mathbb{UHG}$Eval. A concise 20-line demonstration is provided to expedite user initiation, complemented by run.py capable of initiating experiments via the command line.
@@
III-D Experimental Setup
To establish a robust experimental framework, our configuration includes prompt engineering, ensuring equilibrium between positive and negative examples, optimizing hyper-parameters, and configuring evaluators.
Prompt engineering. The prompt engineering technique employed is “intent + instruction + 3-shot (explainable) prompting.” Intent delineates the LLM’s role, instruction outlines the task for the LLM to execute, and the prompt incorporates three examples to aid the LLM’s few-shot learning [1]. Furthermore, political content in examples is prohibited to adhere to content policies from model service providers. Explainable prompting entails not merely acquiring results but also eliciting the model’s rationale behind its responses, regardless of the impact on evaluation speed and cost. In discriminative and selective evaluations, it is indiscernible whether the model is conjecturing the outcome or discerning the presence of hallucinations. Consequently, the use of explainable prompting enables the validation of the model’s confidence through the analysis of experimental results.
Balancing positive and negative examples. To guarantee the reliability of experimental outcomes for all LLMs, we meticulously balance examples in discriminative and selective evaluations. Specifically, the LLM under evaluation will encounter an equal number of examples with and without hallucinations. This approach addresses the tendency of some models to learn patterns from the three examples in the prompts and produce conjectural rather than reasoned responses when making judgments. Such a tendency can introduce a considerable bias towards certain outcomes. An imbalance could complicate the analysis of experimental outcomes.
Hyperparameter settings. Managing parameters for heterogeneous LLMs is a multifaceted endeavor, as different LLMs feature unique interface designs, and the same parameters can have varying implications across LLMs. For example, the level of determinism influenced by the temperature parameter varies. Despite these challenges, we commit to the principle of “guaranteeing overall output determinism while allowing for slight randomness, and aiming for consistent parameter settings across models.” Consequently, we configured parameters including temperature, top_p, top_k [1], and random seed. To ensure output determinism and improve reproducibility, we set the temperature to 0.1. Considering that OpenAI models advise against adjusting temperature and top_p simultaneously, we minimally altered top_p, setting it at 0.9. We set top_k to 5, which is effective for certain models. To further enhance reproducibility, we established a seed for random number generators, setting it at 22.
Evaluator Settings. Discriminative evaluation encompasses assessments at two levels of granularity: sentence-level and keyword-level. Prompt design for both levels utilizes the “intent + instruction + 3-shot (explainable) prompting” approach. Furthermore, we maintain a balanced representation of positive and negative examples at both levels. For discriminative evaluation, accuracy serves as the metric. Selective evaluation adheres to the identical prompt design. Each evaluated LLM is presented with one positive and one negative example for every news item. To uphold the integrity of the evaluation, the order of positive and negative examples is randomly alternated with a 50% chance. Accuracy is also employed as the evaluation metric. The generative evaluation’s prompt design adheres to the principle of $\mathbb{UHG}$. Evaluation metrics comprise 4-gram BLEU (BLEU-4), longest common subsequence-based ROUGE (ROUGE-L), kwPrec, and BERTScore.
III-E Results and Analysis
Results are presented in Table IV, Table V, and Table VI.
TABLE IV: Discriminative (Keyword and Sentence Level) and Selective Evaluation Results
@@ III-E Results and Analysis
  • Note: In the table, #kws denotes the number of keywords and #valid denotes number of valid evaluations. In the same column of values, optimal values are bolded and suboptimal values are underlined. The same as below.
TABLE V: Generative evaluation results
@@
TABLE VI: Evaluation Results by Different Types
KNO DOC GEN NUM
@@
  • Note: Read by row. In the same row of values, optimal values are bolded and suboptimal values are underlined.
@@
Discriminative evaluation. Initially, the GPT series models’ performance is notably superior. In the keyword-level assessment, GPT4-0613 and GPT3.5-Turbo respectively achieve the top two rankings. At the sentence level, GPT4-0613 and GPT4-1106 respectively attain the first and second spots. As previously hypothesized, discriminative evaluation requires robust foundational capabilities from LLMs, such as knowledge recall, utilization, and judgment. The GPT series models markedly surpass other models, showcasing their formidable foundational capabilities. Moreover, a comparison of experimental outcomes at the keyword and sentence levels reveals that accuracy is generally superior at the keyword level. This could stem from the fact that the hallucinated continuations in our dataset exhibit sufficient fluency, aligning with the fluency distribution of LLM outputs. This can potentially confuse the evaluated LLM, complicating the judgment of the continuation’s authenticity. Conversely, keywords bypass fluency concerns, rendering keyword-level evaluation more amenable to LLMs. This observation implies that detecting hallucinations could be more dependable at the keyword level compared to the sentence level.
@@
Selective evaluation. Firstly, GPT4-1106 clinches the top spot, reaffirming the formidable foundational capabilities of the GPT series models. Concurrently, Xinyu2-70B attains second place, excelling as a model trained on the Chinese news corpus. This achievement, to a degree, confirms the merit of domain-specific LLMs. Secondly, when comparing the outcomes of the selective evaluation with those of the discriminative evaluation at the sentence level, most LLMs exhibit improved accuracy. This is consistent with our prior conjecture that furnishing LLMs with more contrasting information alleviates the demand on the model’s fact recall, thus diminishing the challenge of selective evaluation. Therefore, we posit that selective evaluation is comparatively simpler for LLMs. Thirdly, a decline is observed in discriminative evaluation outcomes from GPT4-0613 to GPT4-1106, whereas selective evaluation outcomes register a notable increase of around 5%. This substantiates the “seesaw phenomenon,” wherein certain capabilities are enhanced while others may regress, in tandem with the model’s upgrade [30]. This suggests that the decision to either enhance a single capability individually or to balance multiple capabilities is critical.
Generative evaluation. Firstly, InternLM-20B secures two top spots, one runner-up position, and boasts the longest average generation length. This reflects the model’s superior credibility in content generation. However, its kwPrec score is modest, indicating potential for enhancement in keyword-level information generation. Secondly, Xinyu2-70B captures one top spot, two runner-up positions, and has the second-longest average generation length, underscoring its strong credibility in content generation. Its sole underperformance is in the ROUGE metric, which is recall-oriented. Conversely, BLEU and kwPrec are precision-oriented, suggesting the model is adept at delivering consistent output yet faces challenges with factual recall. Thirdly, Aquila-34B achieves the pinnacle in kwPrec scoring, signaling a notable edge in generation quality. However, this could be attributed to its comparatively shorter average generation length. kwPrec assesses the coverage of extended tokens (i.e., keywords), allowing for brief continuations with limited keywords to secure higher keyword coverage in relation to reference information. Fourthly, Baichuan2-53B registers a high ROUGE score, indicative of its proficiency in fact recall from the parameters, demonstrating accurate factual retrieval. Fifthly, the GPT series exhibits subpar performance, owing to the insubstantial Chinese data in its training corpus. For example, the Chinese data incorporated in GPT’s training from the Common Crawl corpus comprises less than 5% [^4] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages.html.
@@
Evaluations by Type. Given the categorization of news into four types, we can proceed with an in-depth analysis. We focus on selective evaluation results and perform a comprehensive breakdown analysis of these across the four types, as illustrated in Table VI. Initially, the majority of LLMs demonstrate enhanced accuracy for knowledge-intensive and document-intensive news. This observation is consistent with the general consensus that the training datasets for LLMs typically include substantial human knowledge and official documentation of major historical events. Furthermore, the majority of LLMs show reduced accuracy in general and number-intensive news. General news often contains societal minutiae, which are not the focus of LLM training, potentially leading to a deficiency in this factual domain within the model parameters. Regarding number-intensive news, it poses a considerable challenge for most LLMs, given that encoding identical numbers with varied historical meanings is complex. Lastly, GPT4-1106 attains especially high scores in the demanding number-intensive news, which might be attributed to its sophisticated parameterization for numerical data handling.
III-F Discussion

### v2

@@ 2.4 Automatic Labeling and Human Rechecking
Through hallucination ranking, we can identify continuations that are both articulately expressed and likely to contain hallucinations. To detect continuations with confirmed hallucinations, we propose an annotation scheme that utilizes keywords, which includes automatic labeling and subsequent human verification, as shown in Fig. 4.
Figure 4: Labeling and rechecking. (In Chinese: Fig. 12)
Automatic labeling
@@
We utilize the keywords identified by GPT3.5-Turbo from the candidate continuations, similarly to the process used in the computation of kwPrec previously. These keywords act as the focal points for subsequent verification. Thereafter, we employ GPT4-0613 OpenAI (2023) to perform annotation on these keywords. It evaluates the validity of the keywords in the continuations by conducting a cross-reference with the provided original news and provides explanations for any detected unreasonable keywords.
Human rechecking
@@
We undertake a manual, one-to-one verification process by analyzing the annotated results and explanations provided by GPT4-0613 against the original news. This step ensures the accuracy of the machine-generated annotations. In the end, instances verified as accurate by annotators comprise the final $\mathbb{UHG}$Eval dataset. For details on manual annotation, please refer to Appendix A.1.
2.5 Dataset Statistics
Starting with 17,714 candidate hallucinated continuations, we curated a dataset of 5,141 hallucinated continuations, as detailed in the basic statistics in Table 2. For further analysis of statistics and an example of the dataset, please refer to the Appendix A.2 and the Appendix A.3, respectively.
DOC
KNO
NUM
GEN
#news
1242
320
2431
1148
avg. #hallu. kw.
2.15
1.99
2.54
2.12
avg. #kw.
8.43
8.09
8.07
8.17
#hallu. kw. / #kw.
25.47%
24.61%
31.44%
26.00%
avg. len. contn.
46.77
48.36
44.47
45.97
avg. len. begin.
102.15
102.66
103.20
102.86
avg. len. refer.
634.17
618.90
624.47
632.47
@@ 2.5 Dataset Statistics
Table 2: Dataset basic statistics. # denotes quantity, avg. denotes average, len. denotes length, contn. denotes hallucinated continuations, begin. denotes news beginnings, and refer. denotes reference information.
3 Experiments
3.1 Models
@@
Given that our dataset is tailored for the Chinese language generation domain, we selected eight widely used Chinese LLMs and three models from OpenAI. These LLMs are from eight base models: Aquila2 Base BAAI (2023), Baichuan2 Base Yang et al. (2023a), GLM Base Du et al. (2022), GPT Base [^2] https://openai.com, InternLM Base InternLM (2023), Qwen Base Bai et al. (2023), BLOOMZ Base Muennighoff et al. (2023), and LLaMA2 Base Touvron et al. (2023). Refer to the Appendix B.1 for a detailed overview of the LLMs used in the experiments.
3.2 Evaluation Forms
In this study, we conducted a detailed analysis of evaluation methods across three dimensions: form, metric, and granularity. A more comprehensive report can be found in the Appendix B.2. Here, we introduce the three forms of evaluation.
Firstly, there is the discriminative evaluation, which involves having the model determine whether a continuation contains hallucinations. Secondly, similar to discriminative evaluation, selective evaluation allows LLMs to choose the continuation without hallucinations from options with and without such content. Lastly, we have generative evaluation. Specifically, the LLM under evaluation is provided with a beginning text and is then tasked with generating a continuation. Subsequently, various reference-based techniques are employed to assess whether the generated continuation includes hallucinations.
3.3 Evaluation Framework
@@
To accommodate different forms of evaluation methods, we have developed a data-secure, easy-to-extend, and easy-to-use evaluation framework, as illustrated in Fig. 5. Refer to Appendix B.3 for a more detailed understanding of the various layers of the framework.
@@
3.4 Experimental Setup
To establish a robust experimental framework, we have set up some configurations as follows.
Prompt engineering
We apply the technique of “intent + instruction + 3-shot (explainable) prompting.” Intent delineates the role, instruction outlines the task, and the prompt incorporates three examples to aid the few-shot learning Zhao et al. (2023). Furthermore, political content in examples is prohibited to adhere to content policies from model service providers. Explainable prompting entails not merely acquiring results, but also eliciting the model’s rationale behind its responses. Refer to Appendix E to view the complete prompt templates.
Example Balancing
To guarantee the reliability of experimental outcomes for all LLMs, we meticulously balance examples in discriminative and also in selective evaluations. Specifically, the LLM under evaluation will encounter an equal number of examples with and without hallucinations.
Hyperparameter settings
Managing parameters for heterogeneous LLMs is a multifaceted endeavor, as different LLMs feature unique interface designs, and the same parameters can have varying implications across LLMs. Despite these challenges, we commit to the principle of “guaranteeing overall output determinism while allowing for slight randomness, and aiming for consistent parameter settings across models.” Consequently, we set the temperature to 0.1, the top_p to 0.9, the top_k to 5, and the random seed to 22.
Metrics
For discriminative and selective evaluation, accuracy serves as the metric. For generative evaluation, metrics consist of 4-gram BLEU (BLEU-4), the longest common subsequence-based ROUGE (ROUGE-L), kwPrec, and BERTScore.
3.5 Results and Analysis
Results are presented in Table 3 and Table 4.
@@ 3.5 Results and Analysis
Generative
@@
Table 3: Discriminative, selective, and generative evaluation results. #kws denotes the number of keywords and #valid denotes the number of valid evaluations. In the same column, optimal values are bolded, and suboptimal values are underlined.
KNO
DOC
GEN
NUM
Aquila-34B
59.55%
@@
Table 4: Evaluation by different types. In the same row, optimal values are bolded, and suboptimal values are underlined.
Discriminative evaluation
@@
Initially, the GPT series models’ performance is notably superior in discriminative evaluation, showcasing their formidable foundational capabilities in knowledge recall, utilization, and judgment. Moreover, a comparison of experimental outcomes at the keyword and sentence levels reveals that accuracy is generally superior at the keyword level. This could stem from the fact that the hallucinated continuations in our dataset exhibit sufficient fluency, aligning with the fluency distribution of LLM outputs. This can potentially confuse the evaluated LLM, complicating the judgment of the continuation’s authenticity. Conversely, keywords bypass fluency concerns, rendering keyword-level evaluation more amenable to LLMs. This observation implies that detecting hallucinations could be more dependable at the keyword level compared to the sentence level.
Selective evaluation
@@
Firstly, GPT4-1106 clinches the top spot, reaffirming the formidable foundational capabilities of the GPT series models. Concurrently, Xinyu2-70B attains second place, excelling as a model trained on the Chinese news corpus. This achievement, to a degree, confirms the merit of domain-specific LLMs. Secondly, when comparing the outcomes of the selective evaluation with those of the discriminative evaluation at the sentence level, most LLMs exhibit improved accuracy. We think, furnishing LLMs with more contrasting information alleviates the demand for the model’s fact recall, thus diminishing the challenge of selective evaluation. Therefore, we posit that selective evaluation is comparatively simpler for LLMs.
Generative evaluation
Overall, InternLM-20B, Xinyu2-70B, and Aquila-34B have achieved commendable results, but the performance of Aquila-34B could be attributed to its comparatively shorter average generation length. Additionally, the GPT series exhibits subpar performance, possibly due to the insubstantial amount of Chinese data in its training corpus. After all, the Chinese data incorporated into GPT’s training from the Common Crawl corpus comprises less than 5% [^3] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages.html.
Evaluations by Type
@@
We focus on selective evaluation results and perform a comprehensive breakdown analysis of these across the four types, as illustrated in Table 4. Initially, most LLMs demonstrate enhanced accuracy for knowledge-intensive and document-intensive news. This may be because the training datasets for LLMs typically include substantial human knowledge and official documentation of major historical events. Furthermore, the majority of LLMs show reduced accuracy in general and number-intensive news. General news often contains societal minutiae, which are not the focus of LLM training. Regarding number-intensive news, it poses a considerable challenge for LLMs, given that encoding identical numbers with varied historical meanings is complex. However, GPT4-1106 attains especially high scores in the demanding number-intensive news.
3.6 Further Discussion